Circular No. 8291

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

2004 DW A slow-moving body reported by M. E. Brown, D. L. Rabinowitz, and

C. A. Trujillo, on CCD images obtained with the 1.2-m Schmidt telescope at Palomar on Feb. 17 (cf. MPECs 2004-D09 and 2004-D13), is apparently the second-largest known transneptunian object (after Pluto), with an absolute magnitude H = 2.2 (assumed albedo parameter G = 0.15). Subsequent observations have been found on CCD images and sky-survey photographs back to 1951 by R. Matson, M. Meyer, and R. Stoss, leading to plutino-type orbital elements (a = 39.5 AU, e = 0.22, i = 20^{o}.6; cf. MPEC 2004-D15).

COMETS C/2003 F5, G4, H8-H12, AND J2-J7 (SOHO) Further to IAUC 8283, measurements have been reported by K. Bat-

tams (for C/2003 H12 and C/2003 J6) and by K. McGleam (remaining comets) for numerous Kreutz sungrazing comets that were found on SOHO C2 website images by R. Kracht (C/2003 F5, C/2003 H11, C/2003 J3, and C/2003 J6), X. Leprette (C/2003 G4), X.-m. Zhou (C/2003 H8 and C/2003 H9), T. Hoffman (C/2003 H10), R. Matson (C/2003 H12, C/2003 J2, and C/2003 J4), P. Shkreby (C/2003 J5), J. Sachs (C/2003 J6), and H. Otterstedt (C/2003 J7). C/2003 H12 was noted as being very faint and visible on only four images. C/2003 F5, C/2003 J6, and C/2003 J7 were also visible in C3 coronagraph images. The reduced astrometry and orbital elements by B. G. Marsden appear on the MPECs cited below.

Comet  2003 UT  α_{2000}  δ_{2000}  MPEC
C/2003 F5  Mar. 29.238  0^{h}59.^{m}0  + 1^{o}39′  2004-C53
C/2003 G4 Apr. 2.438  0 53.2  + 4 21  2004-C53
C/2003 H8  22.983  2 08.4  +11 23  2004-C53
C/2003 H9  24.643  2 14.6  +11 54  2004-C53
C/2003 H10  28.976  2 30.2  +13 11  2004-C54
C/2003 H11  29.518  2 29.1  +13 02  2004-C54
C/2003 H12  30.868  2 36.9  +13 42  2004-C54
C/2003 J2  May 1.226  2 38.2  +13 51  2004-C54
C/2003 J3  5.785  2 55.1  +15 07  2004-D08
C/2003 J4  5.910  2 56.1  +15 12  2004-D08
C/2003 J5  8.288  3 04.8  +15 46  2004-D08
C/2003 J6  9.613  3 12.1  +15 47  2004-D08
C/2003 J7  14.738  3 30.5  +16 03  2004-D08
2004 February 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green